# Embedded coherent structures from MHD to sub-ion scales in turbulent solar wind at 0.17 AU

### v2

@@ 8 Conclusion and discussions
We thank the reviewer for his/her careful reading and analysis of our paper. The comments were very helpful to improve the paper. AV acknowledges funding support from the Initiative Physique des Infinis (IPI), a research training program of the Idex SUPER at Sorbonne Université. OA acknowledges funding support from CNES.
@@ A.1.1 Monopole Alfvén vortex
A monopole Alfvén vortex is localized within the cylinder of the radius $a$, and the axis of the cylinder is aligned with $\mathbf{B}_{0}$. The model assumes that the total current inside $r<a$ is zero. If $\delta\mathbf{B}_{\perp}$ is continuous at $r=a$, it implies the condition $J_{1}(ka)=0$, where $J_{1}$ is the first order Bessel function. This defines the parameter $k$ for a given radius $a$. The monopole vortex solution writes (in dimentionless units, see Petviashvili & Pokhotelov (1992)):
@@ A.1.2 Dipole Alfvén vortex
As in case of the monopole vortex, the dipole vortex is a coherent structure localised inside the cylinder of the radius $a$ and the generalised Alfvén relation $\delta\mathbf{V}_{\perp}/V_{A}=\xi\,\delta\mathbf{B}_{\perp}/B_{0}$ is assumed.
The particular property of the dipole Alfvén vortex model is that it’s axis can be inclined by a small angle $\theta$ with respect to the background magnetic field $B_{z}=B_{0}$. We define $\alpha=\tan(\theta)$. Without restriction of generality, let the axis of the vortex be in the $(y,z)$-plane. If $\theta\neq 0$, the dipole vortex propagate along $y$ with the speed $u\propto\alpha$. The continuity of $\delta\mathbf{B}_{\perp}$ at $r=a$ requires that the amplitude of the dipole vortex is not arbitrary, but defined by $\alpha$ and $k$.
In the reference frame moving with the vortex, the dimentionless vector potential of the dipole vortex is (Petviashvili & Pokhotelov, 1992; Alexandrova, 2008):
@@ A.2 Current sheets
MHD classification of current sheets include rotational (RDs) and tangential (TDs) discontinuities (e.g., Baumjohann & Treumann, 1997; Tsurutani et al., 2011). A typical method to distinguish RD from TD is based on the normalised change in magnetic field magnitude $\Delta B/B$ across the discontinuity (which is zero for RD) and the normal magnetic field component $B_{n}/B$ (which is zero for TD). However, observations showed that current sheets can combine both properties of RTs and TDs (e.g., Neugebauer, 2006; Artemyev et al., 2019).
@@ A.2.1 Rotational discontinuity
RDs are characterised by the correlated rotation of magnetic field and velocity (Walen relation in case of the pressure isotropy: $\mathbf{\delta B}/B_{0}=\pm\mathbf{\delta V}/V_{A}$), constant magnetic and plasma pressures across the sheet ($\Delta B/B=\Delta P/P=0$). Plasma on the both sides of a RD is magnetically connected, i.e., $B_{n}\neq 0$.
@@
We select $\Delta\zeta=120^{\circ}$ as in the example that we discuss in Section B.3. According to the statistical study of the current sheets from the first PSP perihelion, thin current sheets have smaller magnetic field rotation angles $\Delta\zeta\sim 21^{\circ}\times(\ell/d_{i})^{0.32}$, where $\ell$ is the CS thickness and $d_{i}$ is the proton inertial length (Lotekar et al., 2022). For rotational discontinuities with smaller $\Delta\zeta$ the polarization becomes closer to linear, i.e., closer to case of tangential discontinuity model discussed in the Section A.2.2. In terms of eigenvalues, $\lambda_{3}/\lambda_{2}$ increases while $\Delta\zeta$ decreases. The selection of $\Delta\zeta=120^{\circ}$ corresponds to a high-amplitude RD. The other ones cannot be distinguished from TD with the polarisation and MVA eigenvalue ratios.
@@ A.2.2 Tangential discontinuity
TDs separate two magnetically disconnected regions of plasma, so normal magnetic field component is zero, $B_{x}=B_{n}=0$. We use the Harris-like current sheet model, with a constant guide field $B_{z}=B_{0}$ (Harris (1962)):
@@
In presence of the strong guide field $B_{0}\gg B_{t}$, the current density is quasi-parallel to magnetic field. So, the current sheet is quasi-force-free in accordance with what was found in observations (Artemyev et al., 2019).
@@ A.3 Magnetic holes
Magnetic hole represents a localized magnetic field modulus decrease. MHD-scale magnetic holes (with crossection widths ranging from $\sim 10\,\rho_{p}$ to $\sim 10^{3}\rho_{p}$, where $\rho_{p}$ is the proton Larmor radius), are quite rare events: at 1 AU the occurrence rate of 0.6 per day was observed by Stevens & Kasper (2007). Closer to the Sun the occurrence rate was found to be higher: from 2.4 per day at 0.7 AU to 3.4 per day at 0.3 AU (Volwerk et al., 2020).
@@ B.1 Example 1
Figure 11: Example 1: A tangential discontinuity on MHD scales. The four panels on the left represent magnetic fluctuations near the central time $t_{0}$ = 0:36:27 UT. (a) Original magnetic field in RTN reference frame, (b) bandpass filtered magnetic fluctuations at MHD, (c) ion and (d) sub-ion frequency ranges, see equations 6 and 7 in local MVA reference frame. (e-h) Corresponding polarisation plots. The orientation of MVA basis vectors $\{\mathbf{e}_{i}\}_{i=1,2,3}$, with respect to the background magnetic field $\mathbf{B}_{0}$ is provided in the legend with $\theta_{B_{0},i}$ angles. Angles for well-defined MVA basis vectors are shown in bold. Panels of the right column show different plasma parameters at MHD scales: (i-k) the correlation between $(\mathbf{B}-\langle\mathbf{B}\rangle)/B_{0}$ and $(\mathbf{V}-\langle\mathbf{V}\rangle)/V_{A}$, (l) the magnetic field and velocity modulus ($|B|$ and $|V|$), (m) the electron density $N_{e}$, (n) proton and electron temperatures $T_{p}$, $T_{e,total}$.
@@
The first event was observed on November 6, 2018, at $t_{0}=$00:36:27 UT. Figure 11(a) shows magnetic field data in the RTN reference frame during 200 s around $t_{0}$. Here, $B_{T}$ and $B_{N}$ components change sign in the center, magnetic field rotates by the angle $\Delta\zeta=80^{\circ}$. Thus, this is an example of a current sheet. PSP crosses this current sheet in $\Delta t\simeq 1.5$ s, so its thickness is about $V\Delta t\sim 450$ km. The flow-to-field angle is $\Theta_{BV}=113^{\circ}$, so the PSP cross this structure under a quasi-perpendicular angle. The polarization of these fluctuations in the plane $(B_{N},B_{T})$ is shown in panel (e). The out of plane $B_{R}$ is negative and nearly constant during the considered time interval, so this discontinuity is not at the edge of a switchback.
@@
Figure 11(b) shows band-pass filtered MHD inertial range magnetic fluctuations $\delta\mathbf{B}_{MHD}$ during the same 200 s around $t_{0}$ in the MVA reference frame. The grey horizontal bands indicates ±$\pm 2\sigma_{MHD}$ (two standard deviations of the random phase signal at MHD scales). The discontinuity in the center is due to the presence of the current sheet detected already in the raw data with the amplitude $\delta B/B\simeq 1.4$. The shape of $\delta B_{1}$ (red line) around $t_{0}$ is due to the band-pass filtering or a current sheet shown above in panel (a). The corresponding polarization (panel (f)) is nearly linear. In the legend, we indicate the angles between the corresponding mean field $\mathbf{B}_{0}$ and the MVA basis. The MVA basis vectors $\mathbf{e}_{1}$, $\mathbf{e}_{2}$ and $\mathbf{e}_{3}$ are well-defined if both eigenvalue ratios are small, $\lambda_{2}/\lambda_{1}<0.3$ and $\lambda_{3}/\lambda_{2}<0.3$ (Paschmann & Daly, 1998). If only the first (second) of the ratios is below $0.3$, then only $\mathbf{e}_{1}$ ($\mathbf{e}_{3}$) is unambiguously defined. The angles with eigenvalue ratios below $0.3$ are shown in bold in the legend of the polarizaion plane. So, one can see a linear polarization, with the maximum variance direction $\mathbf{e_{1}}$ quasi-perpendicular to $\mathbf{B}_{0}$, $\Theta_{B,1}=80^{\circ}$. The intermediate $\mathbf{e}_{2}$ and minimum $\mathbf{e}_{3}$ variance directions are ill defined.
@@
Figure 11(c) shows a zoom to the time interval of ±$\pm 1$ s around the same central time $t_{0}$. The grey horizontal band indicates ±$\pm 2\sigma_{ion}$. For ion scales, the amplitude is significant, $\delta\mathbf{B}_{ion}/B_{0}\sim 0.2$, but smaller than one. The shape of $\delta\mathbf{B}_{ion}$ is not the filtering remnant of the current sheet, as shown in the Appendix C. The black dashed line shows the fluctuations of magnetic field modulus $\delta|B|$, which are negligible. Here, the local MVA frame is well defined and minimal and maximal variations are perpendicular to the field. The elliptic polarization and the shape of magnetic fluctuations at ion scales resemble the crossing of a dipole vortex (shown in the Figure 6(b)). Thus, we observe a vortex like structure at ion scales embedded in the current sheet at MHD scales.
@@
Figure 11(d) shows the zoom-in to the time interval of ±$\pm 0.07$ s around $t_{0}$. The grey horizontal band indicates ±$\pm 2\sigma_{sub-ion}$. Here magnetic fluctuations $\delta\mathbf{B}_{sub-ion}$ are well localized in time and has a non-negligible fluctuations of the modulus of magnetic field $\delta|B|/\delta B_{1}=0.4$ (black dashed line). Observation of an important compressible component in $\delta\mathbf{B}_{sub-ion}$ is in agreement with a statistical increase of compressibility at sub-ion scale (see the spectrum of compressible fluctuations, Figure 2(b)). Sub-ion scale fluctuations have elliptic polarisation, see panel (h). The maximum MVA eigenvector is quasi-perpendicular to the background magnetic field $\theta_{B_{0},1}=69^{\circ}$. The properties mentioned above can be explained as a compressible vortex (Jovanović et al., 2020), crossed through its center.
@@
Figure 11(i-k) show the magnetic field fluctuations normalised by the background magnetic field $(\mathbf{B}-\langle\mathbf{B}\rangle)/B_{0}$, where $B_{0}=\langle|B|\rangle_{t-t_{0}\in(-100,100)s}=100$ nT, and the proton velocity fluctuations $(\mathbf{V}-\langle\mathbf{V}\rangle)/V_{A}$ normalised by the average Alfvén velocity $V_{A}=B_{0}/\sqrt{\mu_{0}N_{e}m_{p}}=104$ km/s. Both $(\mathbf{B}-\langle\mathbf{B}\rangle)/B_{0}$ and $(\mathbf{V}-\langle\mathbf{V}\rangle)/V_{A}$ are shown in magnetic field MVA reference frame calculated using $(\mathbf{B}-\langle\mathbf{B}\rangle)$ vector over the 200 seconds shown. Magnetic field and velocity variations across the sheet ($\Delta$) are nearly aligned $\alpha(\Delta\mathbf{B},\Delta\mathbf{V})=21^{\circ}$. Variations in $(\mathbf{B}-\langle\mathbf{B}\rangle)/B_{0}$ and $(\mathbf{V}-\langle\mathbf{V}\rangle)/V_{A}$ correlate, but the amplitudes are different $(|\Delta\mathbf{V}|/V_{A})_{1}=0.4\cdot(|\Delta\mathbf{B}|/B_{0})_{1}$.
Thus, the discontinuity does not fulfill the Walen relation $\Delta\mathbf{V}=\pm\Delta\mathbf{B}/\sqrt{4\pi\rho}$ for rotational discontinuities. In presence of pressure anisotropy, the density can change across the discontinuity, and the Walen relation is modified, as follows (Hudson, 1970; Neugebauer, 2006):

### v4

@@ 8 Conclusion and discussions
We thank the reviewer for his/her careful reading and analysis of our paper. The comments were very helpful to improve the paper. AV acknowledges funding support from the Initiative Physique des Infinis (IPI), a research training program of the Idex SUPER at Sorbonne Université. OA acknowledges funding support from CNES. Wavelet software was provided by C. Torrence and G. Compo, and is available at URL: http://atoc.colorado.edu/research/wavelets/.
@@ A.1.1 Monopole Alfvén vortex
A monopole Alfvén vortex is localized within the cylinder of the radius $a$, and the axis of the cylinder is aligned with $\mathbf{B}_{0}$. The model assumes that the total current inside $r<a$ is zero. If $\delta\mathbf{B}_{\perp}$ is continuous at $r=a$, it implies the condition $J_{1}(ka)=0$, where $J_{1}$ is the first order Bessel function. This defines the parameter $k$ for a given radius $a$. The monopole vortex solution writes (in dimensionless units, see Petviashvili & Pokhotelov (1992)):
@@ A.1.2 Dipole Alfvén vortex
As in the case of the monopole vortex, the dipole vortex is a coherent structure localized inside the cylinder of the radius $a$ and the generalized Alfvén relation $\delta\mathbf{V}_{\perp}/V_{A}=\xi\,\delta\mathbf{B}_{\perp}/B_{0}$ is assumed.
The particular property of the dipole Alfvén vortex model is that its axis can be inclined by a small angle $\theta$ with respect to the background magnetic field $B_{z}=B_{0}$. We define $\alpha=\tan(\theta)$. Without restriction of generality, let the axis of the vortex be in the $(y,z)$-plane. If $\theta\neq 0$, the dipole vortex propagates along $y$ with the speed $u\propto\alpha$. The continuity of $\delta\mathbf{B}_{\perp}$ at $r=a$ requires that the amplitude of the dipole vortex is not arbitrary, but defined by $\alpha$ and $k$.
In the reference frame moving with the vortex, the dimensionless vector potential of the dipole vortex is (Petviashvili & Pokhotelov, 1992; Alexandrova, 2008):
@@ A.2 Current sheets
MHD classification of current sheets includes rotational (RDs) and tangential (TDs) discontinuities (e.g., Baumjohann & Treumann, 1997; Tsurutani et al., 2011). A typical method to distinguish RD from TD is based on the normalized change in magnetic field magnitude $\Delta B/B$ across the discontinuity (which is zero for RD) and the normal magnetic field component $B_{n}/B$ (which is zero for TD). To test the criterion $B_{n}/B=0$, the normal to the sheet must be accurately determined. Current sheets are planar structures, so $B_{n}=const$ (due to the divergence-free magnetic field). Therefore, the direction of the normal ($\mathbf{n}$) can be estimated using MVA, namely $\mathbf{n}=\mathbf{e_{3}}$. However, the error of this method is significant (Horbury et al., 2001; Knetter et al., 2004; Wang et al., 2024). Thus, the verification of the criterion by MVA in application to single-spacecraft data is not accurate enough to classify the current sheets in the solar wind as tangential or rotational. Observations showed that current sheets can combine the properties of RDs and TDs, so classifying them as RDs or TDs might be an oversimplification (e.g., Neugebauer, 2006; Artemyev et al., 2019).
@@ A.2.1 Rotational discontinuity
RDs are characterized by the correlated rotation of magnetic field and velocity (Walen relation in case of the pressure isotropy: $\mathbf{\delta B}/B_{0}=\pm\mathbf{\delta V}/V_{A}$), constant magnetic and plasma pressures across the sheet ($\Delta B/B=\Delta P/P=0$). Plasma on both sides of a RD is magnetically connected, i.e., $B_{n}\neq 0$.
@@
We select $\Delta\zeta=120^{\circ}$ as in the example that we discuss in Section B.3. According to the statistical study of the current sheets from the first PSP perihelion, thin current sheets have smaller magnetic field rotation angles $\Delta\zeta\sim 21^{\circ}\times(\ell/d_{i})^{0.32}$, where $\ell$ is the CS thickness and $d_{i}$ is the proton inertial length (Lotekar et al., 2022). For rotational discontinuities with smaller $\Delta\zeta$ the polarization becomes closer to linear, i.e., closer to the case of the tangential discontinuity model discussed in Section A.2.2. In terms of eigenvalues, $\lambda_{3}/\lambda_{2}$ increases while $\Delta\zeta$ decreases. The selection of $\Delta\zeta=120^{\circ}$ corresponds to a high-amplitude RD. The RDs with small $\Delta\zeta$ cannot be distinguished from TD with the polarization and MVA eigenvalue ratios.
@@ A.2.2 Tangential discontinuity
TDs separate two magnetically disconnected plasma regions, so the normal component of the magnetic field is zero, $B_{x}=B_{n}=0$. We use the Harris-like current sheet model, with a constant guide field $B_{z}=B_{0}$ (Harris (1962)):
@@
In the presence of the strong guide field $B_{0}\gg B_{t}$, the current density is quasi-parallel to the magnetic field. So, the current sheet is quasi-force-free in accordance with what was found in observations (Artemyev et al., 2019).
@@ A.3 Magnetic holes
A magnetic hole represents a localized magnetic field modulus decrease. MHD-scale magnetic holes (with cross-section widths ranging from $\sim 10\,\rho_{p}$ to $\sim 10^{3}\rho_{p}$, where $\rho_{p}$ is the proton Larmor radius), are quite rare events: at 1 AU the occurrence rate of 0.6 per day was observed by Stevens & Kasper (2007). Closer to the Sun the occurrence rate is higher: 2.4 per day at 0.7 AU, 3.4 per day at 0.3 AU (Volwerk et al., 2020).
@@ B.1 Example 1
Figure 11: Example 1: A tangential discontinuity on MHD scales. The four panels on the left represent magnetic fluctuations near the central time $t_{0}$ = 0:36:27 UT. (a) Original magnetic field in the RTN reference frame, (b) bandpass filtered magnetic fluctuations at MHD, (c) ion and (d) sub-ion frequency ranges, see equations 6 and 7 in local MVA reference frame. (e-h) Corresponding polarization plots. The orientation of MVA basis vectors $\{\mathbf{e}_{i}\}_{i=1,2,3}$, with respect to the background magnetic field $\mathbf{B}_{0}$ is provided in the legend with $\theta_{B_{0},i}$ angles. Angles for well-defined MVA basis vectors are shown in bold. Panels of the right column show different plasma parameters at MHD scales: (i-k) the correlation between $(\mathbf{B}-\langle\mathbf{B}\rangle)/B_{0}$ and $(\mathbf{V}-\langle\mathbf{V}\rangle)/V_{A}$, (l) the magnetic field and velocity modulus ($|B|$ and $|V|$), (m) the electron density $N_{e}$, (n) proton and electron temperatures $T_{p}$, $T_{e,total}$.
@@
The first event was observed on November 6, 2018, at $t_{0}=$00:36:27 UT. Figure 11(a) shows magnetic field data in the RTN reference frame during 200 s around $t_{0}$. Here, $B_{T}$ and $B_{N}$ components change the sign in the center, the magnetic field rotates by the angle $\Delta\zeta=80^{\circ}$. Thus, this is an example of a current sheet. PSP crosses this current sheet in $\Delta t\simeq 1.5$ s, so its thickness is about $V\Delta t\sim 450$ km. The flow-to-field angle is $\Theta_{BV}=113^{\circ}$, so the PSP crosses this structure under a quasi-perpendicular angle. The polarization of the fluctuations in the plane $(B_{N},B_{T})$ is shown in panel (e). The out-of-plane $B_{R}$ is negative and nearly constant during the considered time interval, so this discontinuity is not at the edge of a switchback.
@@
Figure 11(b) shows band-pass filtered MHD inertial range magnetic fluctuations $\delta\mathbf{B}_{MHD}$ during the same 200 s around $t_{0}$ in the MVA reference frame. The grey horizontal bands indicate ±$\pm 2\sigma_{MHD}$ (two standard deviations of the random phase signal at MHD scales). The discontinuity in the center is due to the presence of the current sheet detected already in the raw data with the amplitude $\delta B/B\simeq 1.4$. The shape of $\delta B_{1}$ (red line) around $t_{0}$ is due to the band-pass filtering or a current sheet shown above in panel (a). The corresponding polarization (panel (f)) is nearly linear. In the legend, we indicate the angles between the corresponding mean field $\mathbf{B}_{0}$ and the MVA basis. The MVA basis vectors $\mathbf{e}_{1}$, $\mathbf{e}_{2}$ and $\mathbf{e}_{3}$ are well-defined if both eigenvalue ratios are small, $\lambda_{2}/\lambda_{1}<0.3$ and $\lambda_{3}/\lambda_{2}<0.3$ (Paschmann & Daly, 1998). If only the first (second) of the ratios is below $0.3$, then only $\mathbf{e}_{1}$ ($\mathbf{e}_{3}$) is unambiguously defined. The angles with eigenvalue ratios below $0.3$ are shown in bold in the legend of the polarization plane. So, one can see a linear polarization, with the maximum variance direction $\mathbf{e_{1}}$ quasi-perpendicular to $\mathbf{B}_{0}$, $\Theta_{B,1}=80^{\circ}$. The intermediate $\mathbf{e}_{2}$ and minimum $\mathbf{e}_{3}$ variance directions are ill-defined.
@@
Figure 11(c) shows a zoom to the time interval of ±$\pm 1$ s around the same central time $t_{0}$. The grey horizontal band indicates ±$\pm 2\sigma_{ion}$. For ion scales, the amplitude is significant, $\delta\mathbf{B}_{ion}/B_{0}\sim 0.2$. The shape of $\delta\mathbf{B}_{ion}$ is not the filtering remnant of the current sheet, as shown in the Appendix C. The black dashed line shows the fluctuations of magnetic field modulus $\delta|\mathbf{B}|$, which are negligible. Here, the local MVA frame is well defined. The minimum and maximum variation directions are perpendicular to the magnetic field. The elliptic polarization and the shape of magnetic fluctuations at ion scales resemble the crossing of a dipole vortex (shown in Figure 6(b)). Thus, we observe a vortex-like structure at ion scales embedded in the current sheet at MHD scales.
@@
Figure 11(d) shows the zoom-in to the time interval of ±$\pm 0.07$ s around $t_{0}$. High-amplitude fluctuations $\delta\mathbf{B}_{sub-ion}$ with respect to the noise level ±$\pm 2\sigma_{sub-ion}$ (grey band) are well localized in time. The modulus of the magnetic field is fluctuating with a significant amplitude $\delta|B|/\delta B_{1}=0.4$ (black dashed line), so the fluctuations are compressible. This is in agreement with a statistical increase of compressibility at the sub-ion range (see the spectrum of compressible fluctuations, Figure 2(b)). The polarization is elliptic, see panel (h). The maximum MVA eigenvector is quasi-perpendicular to the background magnetic field $\theta_{B_{0},1}=69^{\circ}$. These properties can be explained as the crossing of the compressible vortex through its center (Jovanović et al., 2020).
@@
Figure 11(i-k) show the magnetic field fluctuations normalized by the background magnetic field $(\mathbf{B}-\langle\mathbf{B}\rangle)/B_{0}$, where $B_{0}=\langle|B|\rangle_{t-t_{0}\in(-100,100)s}=100$ nT, and the proton velocity fluctuations $(\mathbf{V}-\langle\mathbf{V}\rangle)/V_{A}$ normalized by the average Alfvén velocity $V_{A}=B_{0}/\sqrt{\mu_{0}N_{e}m_{p}}=104$ km/s. Both $(\mathbf{B}-\langle\mathbf{B}\rangle)/B_{0}$ and $(\mathbf{V}-\langle\mathbf{V}\rangle)/V_{A}$ are shown in magnetic field MVA reference frame calculated using $(\mathbf{B}-\langle\mathbf{B}\rangle)$ vector over the 200 seconds shown. Magnetic field and velocity variations across the sheet ($\Delta$) make the angle of $\alpha(\Delta\mathbf{B},\Delta\mathbf{V})=21^{\circ}$. Variations in $(\mathbf{B}-\langle\mathbf{B}\rangle)/B_{0}$ and $(\mathbf{V}-\langle\mathbf{V}\rangle)/V_{A}$ correlate, but the amplitudes are different $(|\Delta\mathbf{V}|/V_{A})_{1}=0.4\cdot(|\Delta\mathbf{B}|/B_{0})_{1}$. Thus, the discontinuity does not fulfill the Walen relation $\Delta\mathbf{V}=\pm\Delta\mathbf{B}/\sqrt{4\pi\rho}$ for rotational discontinuities. In the presence of pressure anisotropy, the density can change across the discontinuity, and the Walen relation is modified as follows (Hudson, 1970; Neugebauer, 2006):
@@
In Figure 11(l-n) we see how magnetic field modulus $|\mathbf{B}|$, velocity modulus $|\mathbf{V}|$, electron density $N_{e}$, ion $T_{i}$ and electron $T_{e}$ temperatures change across the structure. Velocity and temperatures stay nearly constant. At the same time, $|\mathbf{B}|$ and $N_{e}$ are anti-correlated: while magnetic field increases by $\Delta|\mathbf{B}|=10$ nT, density decreases by $\Delta N_{e}\sim 50$ cm-3. This is usually observed for convected structures in pressure balance. The observed properties are typical for a tangential discontinuity, where magnetic field and density are not constant across the discontinuity.
Another property that distinguishes between RD and TD is the magnitude of the normal magnetic field component $B_{n}$. The divergence-free condition implies that $B_{n}$ must be constant in the case of planar geometry. So, MVA minimum variance direction should represent normal to the magnetic sheet, and $B_{3}=B_{n}=$ constant. Next, tangential discontinuities have $B_{n}=0$, but in observations $B_{3}\simeq 70$ nT. These results are obtained in the known limits of MVA since the MVA estimation of the normal to the sheet can differ from normal estimated from multi-spacecraft methods (Horbury et al., 2001; Knetter et al., 2004).
So, to summarize, starting from the largest observed scales and up to the end of the inertial range, we observe a current sheet that can be interpreted as a tangential discontinuity (TD). Ion and sub-ion scale substructures are embedded in this discontinuity. Ion scale structure resembles the dipole Alfvén vortex model (see Section 5 and Figure 11 column (b)). Sub-ion scale structure might represent a compressible vortex (Jovanović et al., 2020). A sketch describing this event is given in Figure 10(a).
B.2 Example 2
The second example is shown in Figure 12 in the same format as the first event in Figure 11. The central time of the event is 01:19:20 UT. In panel (a), the raw magnetic field is shown in the RTN reference frame.
Panel (e) shows the polarization $B_{T}(B_{N})$; out-of-plane $B_{R}$ does not change sign (this structure is not a switchback). The magnetic field deflects twice within the timescale of $\sim 80$ s. The magnetic field before crossing the structure and after are orientated differently: it is rotated by $\Delta\zeta=15^{\circ}$ (see Equation (A3)). This can be due to a weak ($|\Delta\mathbf{B}|/B_{0}=(\mathbf{B}_{t=-100s}-\mathbf{B}_{t=100s})/B_{0}\simeq 0.3$) rotational current sheet, since the ratio of velocity and magnetic field jumps satisfy the Walen relation |$\Delta\mathbf{V}|/V_{A}=1.03\cdot|\Delta\mathbf{B}|/B_{0}$.
Magnetic fluctuations at the MHD scales $\delta\mathbf{B}_{MHD}$ are shown in Figure 12 (b) in the MVA reference frame. The amplitude of the structure $\delta B_{1}\simeq 60$ nT (i.e., from peak to peak $\Delta B/B_{0}\sim 1.2$) well exceeds the level of incoherent signal $222\sigma_{MHD}=$ 22 nT. The direction of the maximum eigenvector $\mathbf{e}_{1}$ is well-distinguished from intermediate ($\mathbf{e}_{2}$) and minimum ($\mathbf{e}_{3}$) directions since $\lambda_{2}/\lambda_{1}=0.06$, and it is perpendicular to the background magnetic field $\theta_{B_{0},1}=88^{\circ}$.
The velocity and magnetic field fluctuations are well-correlated $\delta B/B_{0}=\xi\delta V/V_{A}$, see panels (i-k). The proportionality coefficient is $\xi=0.86$ for magnetic fluctuations at MHD scales and $\xi=0.81$ in the raw data with the mean value subtracted for the same time interval. The black lines in panels (i-k) show the fluctuations of the monopole Alfvén vortex model, along the central crossing. The observed time profiles of fluctuations agree well with the Alfven vortex model. The magnetic field vector before and after the vortex crossing is oriented differently: $|\mathbf{B}(t_{0}-70~{}\text{s})-\mathbf{B}(t_{0}+70~{}\text{s})|/B_{0}=(0.4,0.03,0.03)$ (in the MVA reference frame). Thus, the vortex is embedded in a relatively weak current sheet.
In the plasma rest frame, this vortex propagates with a negligible speed of $\simeq 3$ km/s (see the main text of the paper), so the Taylor hypothesis can be used to estimate its spatial scale. PSP trajectory crosses the structure in a plane nearly perpendicular to $\mathbf{B}_{0}$ ($\theta_{B_{0},V}=71^{\circ}$). The diameter of the vortex can be estimated as $d=\theta_{B_{0},V}\cdot V\Delta t\sim 2.4\cdot 10^{4}$ km. Given that the mean ion inertial length is $d_{i}\simeq 11$ km and the mean ion Larmor radius is $\rho_{i}=5$ km, the vortex diameter $d\simeq 2\cdot 10^{3}d_{i}=5\cdot 10^{3}\rho_{i}$. The variation of the magnetic field modulus is negligible $\delta|\mathbf{B}|/B_{0}=0.03$ as well as the variations of $N_{e}$, $T_{e}$ and $T_{p}$, see panels (l-n). The change of $|\mathbf{V}|$ (grey line in panel (l)) is due to the superposition of the velocity fluctuation of the Alfvén vortex on the bulk solar wind speed.
Figure 12(c) shows the ion scale magnetic fluctuations $\delta\mathbf{B}_{ion}$ located in the center of the MHD scale Alfvén vortex. The maximum amplitude of the fluctuation: $\max(|\delta B_{1}|)=24$ nT, as well as two secondary peaks on the left and right sides exceed well the incoherent threshold ±$\pm 2\sigma_{ion-scale}$ shown in grey. The polarization is elliptical (panel (g)), and the maximum variance is perpendicular to the local field direction. These observed properties are in agreement with an Alfvén vortex crossing with a finite impact distance from its center. The vortex was observed during $\Delta t=1.5$ s. The diameter of the vortex is $d=V\Delta t\sin(\Theta_{BV})=440$ km $=40~{}d_{i}=90~{}\rho_{i}$, where $\Delta t=1.5$ s is the crossing duration, $V=330$ km/s, $\Theta_{BV}=70^{\circ}$, $d_{i}=11$ km, $\rho_{i}=5$ km.